Based on the assumptions of the aggregate expenditure model, fill in the columns for planned investment, government spending, and net exports.Hwk 7 (Ch. 1): Opens 3/21, Due 3/28 6 8 Consider the data presented in the table: investment Actual aggregate expenditure or output (Y) (billions of S) Consumption (C) Planned investmentspending (6) Net exports (Nx) (billions of S) (billions of S) (billions of S) (billions of S) of s) points 399 476 558 630 618 718 eBook 910 Based on the assumptions of the aggregate expenditure model, fill in the columns for planned investment, government spending, and net exports. Instructions: Enter the values into the table above Raferences a. For each level of actual aggregate expenditure, calculate unplanned inventory investment Instructions: Enter the values into the table above?lf the value is negative, then be sure to enter a minus sign. b. What is the equilibrium level of aggregate expenditure in this economy? Instructions: Enter a number rounded to the nearest whole number c. Suppose that planned investment increases by $20 billion. What is the new equilibrium level of aggregate expenditure in this Instructions: Enter a number rounded to the nearest whole number. Prey8 of 8 Next
Answer #1

Aggregate expenditure (AE) = C + Ip + G + NX where

Ip: Planned investment = 120 (autonomous)

G: Government spending = 110 (autonomous)

NX: Net exports = 10 (autonomous)

(a) All values in $ billion

Y C Ip G NX AE Unplanned Inventory Investment = Y - AE
510 310 120 110 10 550 -40
610 390 120 110 10 630 -20
710 470 120 110 10 710 0
810 550 120 110 10 790 20
910 630 120 110 10 870 40

(b) In equilibrium, Y = AE

Equilibrium aggregate expenditure = $710 billion

(c) When Ip rises by $20 billion, new Ip = $140 billion

Y C Ip G NX AE
510 310 140 110 10 570
610 390 140 110 10 650
710 470 140 110 10 730
810 550 140 110 10 810
910 630 140 110 10 890

Equilibrium aggregate expenditure = $810 billion

(d) MPC = Change in C / Change in Y = $(390 - 310) billion / $(610 - 510) billion = $80 billion / $100 billion = 0.80

(e) Multiplier = 1 / (1 - MPC) = 1 / (1 - 0.8) = 1 / 0.2 = 5.00